OpenStudy (anonymous):

The number of algae y in a lake at any time is inversely proportional to the square root of the number x of fish in the lake at any time. Give an equation relating the two values. Please help me guys, missed this one on my test and can't figure it out.

OpenStudy (one098):

\[y=\sqrt{x}\] I think..

OpenStudy (anonymous):

yeah thats what i put and she corrected me with: \[y=\frac{ k }{ \sqrt{x} }\] does that make sense to you?

OpenStudy (one098):

It might, but what does the k stand for?

OpenStudy (anonymous):

Yes that is the correct answer. Inverse means it's in the denominator.

OpenStudy (anonymous):

see thats what im not sure of either, titanic do you know what the k is?

OpenStudy (anonymous):

k is any constant. y times the square root of x always equals k.

OpenStudy (one098):

Ah thats why.

OpenStudy (anonymous):

ooh alright that makes sense now, thanks titanic!

OpenStudy (anonymous):

im reworking all the problems on my test trying to figure them out so i may have some more posts in a bit

OpenStudy (anonymous):

hanks guys

OpenStudy (one098):

good luck @nsellers23

OpenStudy (anonymous):

No problem. And One098 you understand too now right? I can explain further if needed

OpenStudy (one098):

No, I'm good. Thanks. C:

OpenStudy (anonymous):

Ok no problem

OpenStudy (anonymous):

New Problem: Find the inverse of \[f(x) = 2(1.01)^{t}\]

OpenStudy (anonymous):

f(t) not f(x) sorry

OpenStudy (one098):

@nsellers23 what grade are you in?

OpenStudy (anonymous):

y=2(1.01)^t t=2(1.01)^y Solve for y (t/2) = 1.01^y log(t/2) = log(1.01)^y log(t/2) = y*log(1.01) \[y=\frac{\log{\frac{t}{2}}}{\log{1.01}} \]

OpenStudy (anonymous):

I'm a sophmore in college, first time taking calculus for a computer science major.

OpenStudy (one098):

Oh nice.

OpenStudy (anonymous):

thanks i get that one now, i wasnt even close on my original attempt lol

OpenStudy (anonymous):

No problem